

********************************************************************
UARS WEEKLY STATUS REPORT - 30 JAN 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE returned to normal operations last Friday after being in
day-only mode for the yaw maneuver. Otherwise, we continued to operate
in good health gathering solar data on all available orbital tracking
periods. On average, we received 55% of each day's data amounting to
about 400 minutes of solar observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm)
resolution, Mg II, and high resolution monitoring scans
during this past week for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003149 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1993089. MEPS data files between
1993090 and 1993099 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
The MPG attended the 3-TDRSS Operational Readiness Review
meeting held on 29 January 2004.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 January 2004 (Orbit 67570, GMT Day 017) through
23 January 2004 (Orbit 67675, GMT Day 023).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 017/2155 67584 POWER OFF
HRDI 018/1535 67595 POWER OFF
PEM 018/2152 67599 POWER OFF
SUSIM 019/2226 67614 POWER OFF
SOLSTICE 019/2329 67615 FULLTIME TO DAY ONLY
SOLSTICE 023/1326 67669 DAY ONLY TO FULLTIME
SUSIM 023/1614 67671 POWER ON
A Reverse-to-Forward Yaw Around maneuver (number 126) was successfully
executed in Orbit 67626 beginning 020/17:43:39Z and completing at
020/18:20:49Z on 20 January 2004. All instruments were in their off
or standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 55.43 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.40 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.27 and 19.69 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 21
February 2004 (DOY 052).
The beta angle ranged between the angles of 16.8 degrees to 07.2
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 41 degrees on 05 February 2004 (DOY 036).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 0.64 degrees C. Additional
battery performance data for each battery on 16 January and 23 January
2004 are provided in Appendix D.
The clock error ranged from 14.4 to -10.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 08
HALOE = 01
HRDI = 01
PEM = 01
SUSIM = 02
SOLSTICE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week a network anomaly was seen which caused intermittent loss of
telemetry and command capability to the UARS spacecraft. The anomaly
began on the 018/0654 event and did not go away until approximately
018/2011 when the CISCO switches were rebooted. It then reoccurred on
the 021/0233 event and lasted until approximately 021/0500, once again
clearing after a reboot of the CISCO switches. A meeting was called
with IPNOC the following day to discuss the possible cause of this
problem and how to resolve it. IPNOC could not determine the specific
cause of the problem, but requested a few configuration changes that
will help isolate the problem should it reoccur and possibly eliminate
the problem. They also requested that a ping script be installed on
each some of the devices on the network to help in their analysis.
These changes were made and since the changes, the anomaly has not
reoccurred.
Also this week the FOT conducted the end-to-end test of the 3-TDRS
activities. The simulator was brought up and configured to emulate
the spacecraft, 3-TDRS CMS files were FTPed to Eclipse, and the
install and remove procedures were run at time corresponding to actual
realtime events. All subsystems were nominal following each part of
the test and the test was a complete success. This test will be done
one more time on 03 February 2004 with the actual loads that will be
used during the uplink to the spacecraft.
MEETINGS
The FOT attended a Network Anomaly Meeting on 21 January 2004 to
discuss the problems seen in the previous few days and decide on
corrective actions to prevent future anomalies.
The FOT attended the UARS 3 TDRS Status Meeting on 22 January 2004 to
discuss the status of the remaining activities and make sure all
parties were aware of their responsibilities prior to uplink.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 February 2004 (DOY 052).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
The UARS 3-TDRS ORR is scheduled for 29 January 2004 (DOY 029). Also
the 3-TDRS uplink day is scheduled for 04 February 2004 (DOY 035).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 January 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67572 017/0255 04-016 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67584 017/2155 04-017 HALOE POWER OFF
67595 018/1455 04-018 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67595 018/1535 04-019 HRDI INSTRUMENT POWER OFF
67598 018/2011 04-020 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67599 018/2152 04-021 PEM POWER OFF
67611 019/1637 04-022 MPS YAW: MPSCNFG L5 TO L6
67611 019/1711 04-023 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67613 019/2043 04-024 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67614 019/2226 04-025 SUSIM POWER OFF
67615 019/2329 04-026 SOLSTICE FULLTIME TO DAY ONLY OPS
67626 020/1709 04-022 ACS YAW MANEUVER REV TO FWD
67365 021/0749 04-027 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67641 021/1618 04-022 MPS MPSCNFG L6 TO L5 (YAW)
67642 021/1842 04-022 MPS BATSOC (YAW)
67659 022/2125 04-028 SSPP COMP HTR-17 ON
67659 022/2206 04-028 SSPP COMP HTR-17 OFF
67669 023/1326 04-029 SOLSTICE DAY-ONLY TO FULLTIME OPS
67671 023/1614 04-030 SUSIM POWER ON; PART A
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 January 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 January 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67572 017/0259 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67595 018/1500 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
67598 018/2011 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
67611 019/1717 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67613 019/2052 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67635 021/0749 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 January 2004
16 January 2004 - GMT Day 016 Beta = 17.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.36/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +39.2/-22.4 +5.79/+4.50 +29.6/-14.4 25.8 1.03 20.6
3 +11.2/-112.0 +4.50/+3.23 +28.8/-14.4 25.8 1.02 20.0
23 January 2004 - GMT Day 023 Beta = 07.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-44.8 +4.50/+3.87 +28.4/-12.8 26.1 1.03 19.6
3 +11.2/-33.6 +3.23/+2.60 +27.2/-12.8 26.1 1.02 19.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 January 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-004 FOT 67589 018/0624 N/A IP NOCC switch problems
04-006 FOT 67632 021/0233 N/A Eclipse Hung
04-001 FOT 67665 023/0636 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-001 FOT See List Jan 04 30832 Generic Late Acquisition
04-004 FOT 67589 018/0624 N/A IP NOCC switch problems
04-006 FOT 67632 021/0233 N/A Eclipse Hung